# thanks: $0$-0002-5209-872X

### v2

@@ Appendix B CP violation discovery sensitivity
Appendix C Explicit analytic expression for CP violation from disappearance and unitarity
We explicitly demonstrate that it is possible to measure and discover CP violation from disappearance measurements alone. We work in a completely parameterization independent fashion and write: $U_{\alpha i}\equiv|U_{\alpha i}|e^{i\phi_{\alpha i}}$. We also note that CP violating effects are proportional to the Jarlskog invariant Jarlskog (1985)
$J$ $\equiv\Im(U_{\alpha j}^{*}U_{\beta j}U_{\alpha i}U_{\beta i}^{*})$
$=|U_{\alpha j}||U_{\beta j}||U_{\alpha i}||U_{\beta i}|\sin(\phi_{\beta j}+\phi_{\alpha i}-\phi_{\alpha j}-\phi_{\beta i})\,,$ (10)
up to an overall sign, for $\alpha\neq\beta$ and $i\neq j$.
@@ Appendix C Explicit analytic expression for CP violation from disappearance and unitarity
$4|U_{e2}|^{2}|U_{\mu 2}|^{2}|U_{e3}|^{2}|U_{\mu 3}|^{2}\left[1-\sin^{2}(\phi_{e2}-\phi_{\mu 2}-\phi_{e3}+\phi_{\mu 3})\right]\\ =\left(1-|U_{e2}|^{2}-|U_{\mu 2}|^{2}-|U_{e3}|^{2}-|U_{\mu 3}|^{2}\right.\\ \left.+|U_{e2}|^{2}|U_{\mu 3}|^{2}+|U_{e3}|^{2}|U_{\mu 2}|^{2}\right)^{2}\,.$ (17)
Now we take eq. 5 and set $\alpha=e$, $\beta=\mu$, $i=2$, and $j=3$ to get (with the correct sign, which won’t matter in this calculation)
$J=\Im(U_{e3}^{*}U_{\mu 3}U_{e2}U_{\mu 2}^{*})\\ =|U_{e2}||U_{\mu 2}||U_{e3}||U_{\mu 3}|\sin(\phi_{e2}-\phi_{\mu 2}-\phi_{e3}+\phi_{\mu 3})\,.$ (18)
@@
Thus a measurement of only $|U_{e2}|$, $|U_{e3}|$, $|U_{\mu 2}|$, and $|U_{\mu 3}|$, as is possible from performing good $\nu_{e}$ and $\nu_{\mu}$ disappearance experiments, allows for a determination of $J^{2}$. Thus it is possible to show that $J\neq 0$ if nature supports that via disappearance measurements, although it does not indicate if nature prefers neutrinos or antineutrinos since the sign of $J$ is undetermined and must be measured in a CP violating process such as appearance.
Appendix D DUNE event rates and regions of interest
@@
Appendix E Runtime impact
@@
Appendix F Hyper-Kamiokande
@@
Appendix G Tokai to SURF
@@
Experimentally, the low energy of the $\Delta m^{2}_{21}$ minimum would require the beam to be more off-axis than T2K or HK, which further reduces the flux considerably, and the $\nu_{\mu}$ CC cross section experiences considerable suppression from the muon mass threshold making an experiment of this nature extremely unfeasible. An additional issue arises which is one of energy resolution. Given that the $\Delta m^{2}_{31}$ and $\Delta m^{2}_{32}$ oscillations have large amplitude and oscillate $\sim 35$ times faster than the larger $\widehat{\Delta m^{2}_{21}}$ oscillation, the effect due to the differing amplitudes of the $\widehat{\Delta m^{2}_{21}}$ oscillations from $\cos\delta$ will be rapidly averaged out unless unrealistically exceptional energy resolution can be achieved. Thus only the effect described after eq. 28 from eq. 24 could be detectable, not the effect described in eqs. 29-30. Therefore despite its theoretical interest, this determination of $\cos\delta$ is extremely unfeasible, although theoretically possible.

### v3

@@ Appendix A CP violation discovery sensitivity
Appendix B Explicit analytic expression for CP violation from disappearance and unitarity
We explicitly demonstrate that it is possible to measure and discover CP violation from disappearance measurements alone. We work in a completely parameterization independent fashion and write: $U_{\alpha i}\equiv|U_{\alpha i}|e^{i\phi_{\alpha i}}$. We also note that CP violating effects are proportional to the Jarlskog invariant Jarlskog (1985) written in eq. 5.
@@ Appendix B Explicit analytic expression for CP violation from disappearance and unitarity
$4|U_{e2}|^{2}|U_{\mu 2}|^{2}|U_{e3}|^{2}|U_{\mu 3}|^{2}\left[1-\sin^{2}(\phi_{e2}-\phi_{\mu 2}-\phi_{e3}+\phi_{\mu 3})\right]\\ =\left(1-|U_{e2}|^{2}-|U_{\mu 2}|^{2}-|U_{e3}|^{2}-|U_{\mu 3}|^{2}+|U_{e2}|^{2}|U_{\mu 3}|^{2}+|U_{e3}|^{2}|U_{\mu 2}|^{2}\right)^{2}\,.$ (16)
Now we take eq. 5 and set $\alpha=e$, $\beta=\mu$, $i=2$, and $j=3$ to get (with the correct sign, which will not matter in this calculation)
$J=\Im(U_{e3}^{*}U_{\mu 3}U_{e2}U_{\mu 2}^{*})=|U_{e2}||U_{\mu 2}||U_{e3}||U_{\mu 3}|\sin(\phi_{e2}-\phi_{\mu 2}-\phi_{e3}+\phi_{\mu 3})\,.$ (17)
@@
It is also easy to check that if one element is zero, say e.g. $|U_{e3}|=0$, then $J=0$. From eq. 20, this is equivalent to showing that $|U_{e2}|^{2}|U_{\mu 2}|^{2}=|U_{e1}|^{2}|U_{\mu}|^{2}$ which follows from squaring the triangle closure in eq. 12.
Thus a measurement of only $|U_{e2}|$, $|U_{e3}|$, $|U_{\mu 2}|$, and $|U_{\mu 3}|$, as is possible from performing good $\nu_{e}$ and $\nu_{\mu}$ disappearance experiments, allows for a determination of $J^{2}$. Thus it is possible to show that $J\neq 0$ – if nature supports that – via disappearance measurements. This approach does not, however, indicate if nature prefers neutrinos or antineutrinos since the sign of $J$ is undetermined and must be measured in a CP violating process such as appearance.
Appendix C DUNE event rates and regions of interest
@@
Appendix D Runtime impact
@@
Appendix E Hyper-Kamiokande
@@
Appendix F Tokai to SURF
@@
Experimentally, the low energy of the $\Delta m^{2}_{21}$ minimum would require the beam to be more off-axis than T2K or HK, which further reduces the flux considerably, and the $\nu_{\mu}$ CC cross section experiences considerable suppression from the muon mass threshold making an experiment of this nature extremely unfeasible. An additional issue arises which is one of energy resolution. Given that the $\Delta m^{2}_{31}$ and $\Delta m^{2}_{32}$ oscillations have large amplitude and oscillate $\sim 35$ times faster than the larger $\widehat{\Delta m^{2}_{21}}$ oscillation, the effect due to the differing amplitudes of the $\widehat{\Delta m^{2}_{21}}$ oscillations from $\cos\delta$ will be rapidly averaged out unless unrealistically exceptional energy resolution can be achieved. Thus only the effect described after eq. 27 from eq. 23 could be detectable, not the effect described in eqs. 28-29. Therefore despite its theoretical interest, this determination of $\cos\delta$ is extremely unfeasible, although theoretically possible.